# Structural Prognostic Event Modeling for Multimodal Cancer Survival Analysis

### v1

@@ Abstract
The integration of histology images and gene profiles has shown great promise for improving survival prediction in cancer. However, current approaches often struggle to model intra- and inter-modal interactions efficiently and effectively due to the high dimensionality and complexity of the inputs. A major challenge is capturing critical prognostic events that, though few, underlie the complexity of the observed inputs and largely determine patient outcomes. These events—manifested as high-level structural signals such as spatial histologic patterns or pathway co-activations—are typically sparse, patient-specific, and unannotated, making them inherently difficult to uncover. To address this, we propose SlotSPE, a slot-based framework for structural prognostic event modeling. Specifically, inspired by the principle of factorial coding, we compress each patient’s multimodal inputs into compact, modality-specific sets of mutually distinctive slots using slot attention. By leveraging these slot representations as encodings for prognostic events, our framework enables both efficient and effective modeling of complex intra- and inter-modal interactions, while also facilitating seamless incorporation of biological priors that enhance prognostic relevance. Extensive experiments on ten cancer benchmarks show that SlotSPE outperforms existing methods in 8 out of 10 cohorts, achieving an overall improvement of 2.9%. It remains robust under missing genomic data and delivers markedly improved interpretability through structured event decomposition.
@@ 1 Introduction
Cancer survival analysis aims to estimate patient mortality risk or survival time, serving as a key tool for personalized prognosis and precision oncology, as it helps quantify outcomes, identify prognostic factors, and guide individualized treatment decisions (cox2004note; jenkins2005survival; wiegrebe2024deep). Whole-slide images (WSIs) and genomics are two widely-used tools in cancer survival analysis, as they capture complementary prognostic signals: WSIs reveal how cancer manifests morphologically in tissue, while genomics uncovers its underlying molecular drivers. Their integration offers the opportunity for more accurate, robust, and biologically grounded survival prediction, which has in turn sparked growing interest in multimodal learning approaches for survival analysis (chen2021multimodal; xu2023multimodal; jaume2024modeling; zhou2025robust).
@@
Effective and efficient multimodal learning for survival analysis faces a fundamental challenge: although patient outcomes are often driven by a small number of critical events (vogelstein2013cancer; hosseini2024computational), WSI and genomic signals are rather complex and high-dimensional. High-resolution WSIs can exceed pixels (song2023artificial) and genomics involve thousands of genes (gyHorffy2021survival). These events manifest as meaningful structural associations within WSI and genomic signals—e.g., in histopathology, immune activity and tumor aggressiveness (chan2024spatial; liang2025spatial) are reflected in the density and proximity of the pixel organizations of tumor-infiltrating lymphocytes and cancer nests, while in genomics, the tumor-promoting effects emerge from the co-activation of PI3K/AKT and MAPK pathways (glaviano2023pi3k). However, learning to uncover such event-related patterns structures from the input signals is difficult for machines: these high-level associations are typically sparse, patient-specific, and unannotated. Compounding the challenge, the enormous scale of input units—millions of WSI pixels and thousands of genomic features—creates massive self-interactions within each modality, further amplified by the need for cross-modal alignment, rendering direct modeling prohibitively expensive.
@@
Existing attempts, e.g., (ilse2018attention; shao2021transmil; chen2021multimodal; xu2023multimodal; jaume2024modeling), fall short of capturing high-level prognostic associations. The most relevant studies are prototype-based methods (zhang2024prototypical; song2024multimodal), which employ fixed prototype representations and therefore lack adaptability to individual patients. For example, PIBD (zhang2024prototypical) approximates a prototypical distribution for each risk group via information bottleneck (tishby2000information; alemi2016deep). Yet, these prototypes are static-once trained, the same set is applied to all patients. MMP (song2024multimodal) takes a different route by adopting Gaussian mixture model (GMM) (bishop2006pattern) to define prototypes for each WSI. However, the image-wise prototypes are unlearnable (i.e., fixed) during training, making end-to-end optimization suboptimal.
@@
In this paper, we introduce SlotSPE, a slot-based framework for structural prognostic event modeling. Inspired by factorial coding (schmidhuber1992learning_factorial; higgins2017beta; greff2020binding), our method compresses high-dimensional multimodal inputs into a compact set of semantic slots using a slot attention module (locatello2020object), where each slot corresponds to a latent prognostic event. Unlike fixed prototypes, these slots are dynamically instantiated for each patient, allowing the model to capture individualized prognostic patterns while preserving a concise representation. To reduce slot redundancy and encourage slots to capture distinct prognostic events, we introduce selective slot activation through a Mixture-of-Experts–style decoder (hampshire1989connectionist; jacobs1991adaptive; jordan1994hierarchical; shazeer2017outrageously), which activates only the most predictive slots for each patient. Beyond compression, we further model cross-modal interactions by incorporating a biological prior: histopathological phenotypes often reflect underlying molecular events, revealing molecular-morphology mappings (wang2021predicting; liu2025spatial). Guided by this insight, we design a cross-modal reconstruction task where omics-derived slots are encouraged to predict gene expression from histology image. This alignment not only enforces biologically meaningful interactions across modalities but also improves robustness in missing genomic data.
@@ 2 Related Works
Single Modality. The advent of deep learning marks a paradigm shift in cancer prognosis, where artificial neural networks (NN) can uncover prognostic patterns from radiological (aerts2014decoding; lou2019image; xu2019deep; yao2021deepprognosis), pathological (ciresan2012deep; cirecsan2013mitosis), and molecular data (chaudhary2018deep; ching2018cox) that were previously inaccessible to conventional methods. In histopathology, whole-slide images (WSIs) under the multiple instance learning (MIL) (dietterich1997solving; maron1997framework) paradigm—where a slide is represented as a bag of patches with only slide- or patient-level survival labels—have become standard (hou2016patch; ilse2018attention; shao2021transmil; lu2021data). Recent MIL approaches address the challenge of abundant patches by using clustering (e.g. k-means) (yao2020whole; vu2023handcrafted; yang2024scmil; zhou2024cohort), graph models (li2018graph; chen2021whole), and hierarchical WSI representations (shao2023hvtsurv). Meanwhile, genomic methods moved from simple MLP or shallow nets on high-dimensional gene profiles (haykin1998neural; klambauer2017self) to pathway-based representations that group genes into biologically meaningful sets for better interpretability (zhao2021deepomix; hou2023pathexpsurv; wang2024survconvmixer). Building on these advances, our method preserves the MIL standard for histopathology, adopts pathway-level genomic features.
@@
Multiple Modality. Multimodal approaches, especially the integration of histopathology and genomics, have attracted increasing attention because combining morphological and molecular information often yields better prognostic performance than either modality alone (chen2020pathomic; chen2021multimodal; yuan2025pancancer). Typical strategies include late fusion (e.g., concatenation or Kronecker product) (chen2020pathomic; ding2023pathology; volinsky2024prediction; zhao2025decision; luo2025multimodal), and early fusion via co-attention mechanisms operating at the patch-gene level (chen2021multimodal; xu2023multimodal; zhou2023cross; jaume2024modeling; yuan2025pancancer; wang2025histo; zhou2025robust). Since inputs are complex and high-dimensional, recent work has adopted adaptive token selection (zhang2025adamhf) or compact latent representations by compressing WSI and pathway data via prototypes (liu2025adaptive; song2024multimodal; zhang2024prototypical). The former, however, lacks modeling of interpretable structural associations; the latter fails to address sparsity and patient-specific event variation, relying instead on fixed prototypes.
@@
Structural Event Modeling. A key challenge in multimodal survival analysis lies in handling the high-dimensional histology and omics inputs, which not only incurs substantial computational costs but also complicates modeling of intricate intra-/inter-modal interactions. An important insight is that patient outcomes are often driven by a small number of critical events (vogelstein2013cancer). This motivates approaches that aim to uncover the underlying event-related structures within complex inputs. For examples, PIBD (zhang2024prototypical) and MMP (song2024multimodal) attempt to derive structured representations from complex multimodal data. However, relying on fixed prototypes, these methods cannot dynamically capture the sparse and patient-specific structural events that often determine outcomes. From the perspective of machine learning, this challenge can be framed as a problem of factorial coding, a well-established paradigm concerned with decomposing and compressing high-dimensional data into disentangled latent factors (schmidhuber1992learning_factorial; tipping1999probabilistic; higgins2017beta; kim2018disentangling). Building on this connection, our framework advances beyond static prototypes by introducing learnable slots through slot attention (locatello2020object), modeling prognostic events that are selectively activated per patient and aligned via biologically guided cross-modal reconstruction. This design enables a structured decomposition of survival-related events, binding multimodal prognostic signals to interpretable representations.
@@ 3 Method
Survival prediction is the task of estimating a patient risk of an event (e.g., death) given the pathological observations (e.g. WSIs and genomic) of the patient. In this paper, we tackle the problem of learning such hazard predictor from the multi-modal data. Formally, let the hazard prediction function be denoted as , representing the probability of death occurring at time t (i.e., survival up to time t). The model is trained using the negative log-likelihood (NLL) loss (zadeh2020bias), denoted as (see Appendix B.1 for details). Each patient sample is represented as , where indexes the patient, denotes the pathology data, denotes the genomic profile, , and is the censoring indicator ( for an observed event, for right-censoring). Operationally, following standard practices in computational pathology (chen2021multimodal; jaume2024modeling), WSI and genomic inputs are pre-processed by dedicated image and gene encoders and represented as bags of instances: and , where and denote the numbers of WSI patches and biological pathways, respectively, and is the feature dimension. When modeled with state-of-the-art transformers like prior works (xu2023multimodal; zhou2023cross; jaume2024modeling), the resulting intra- and inter-modal interactions scale quadratically as , creating a major challenge for both efficacy and efficiency.
@@ 3.1 Slot-based Information compression
Inspired by factorial coding (schmidhuber1992learning_factorial; higgins2017beta), which aims to disentangle data into independent latent factors, we perform structural decomposition for two bags of histology and genomic instances, i.e., and , using a slot attention module (locatello2020object). This allows us to compress the large input sets into much smaller slot-based representations, significantly reducing computational overhead. These slots are then decoded using a MoE-style decoder that selectively activates patient-specific slots, promoting both sparsity and competition among them.
Slot Attention. Slot attention (locatello2020object) learns to encode distinct structural patterns present in the input into a finite set of latent vectors, i.e., the “slots”. Since WSI and genomic data inherently contain massive redundancy (zhang2024prototypical; song2024multimodal), it is desirable to aggregate them into a smaller set of compact representations while preserving their semantic content. This makes slot attention particularly well-suited for our task, where the resulting compact representations can be interpreted as latent prognostic events modeled by slots.
We re-organize the input features for both modalities as matrices, . With a set of slots, organized as , initialized from some learnable distributions, the slot attention compresses into slots through iterations of cross-attention-based routing. At each iteration , the slots attend to the input to produce updated representations, which are then refined using a multilayer perceptron (MLP). Slot updates across iterations are applied recurrently using a recurrent neural network (RNNs) (hochreiter1997long; cho2014properties) to preserve temporal consistency and iterative refinement. Mathematically, each slot is obtained by:
@@
Then the top- slots are selected using the Gumbel-Top- trick (gumbel1954statistical; maddison2014sampling; kool2019stochastic) with Straight-Through Estimation (ST) (jang2016categorical), producing a differentiable -hot mask (see Appendix B.2 for details). The selected slots are reweighted by masking and renormalizing the softmax weights:
@@ 3.2 Multimodal Slots Interaction with Bilogical Guidance
Cross-modal Reconstruction using Biological Prior. To ensure effective cross-modal interaction, we introduce cross-modal reconstruction (Figure 2(b)) to uncover molecular-morphology mappings reflecting the fact that molecular events give rise to histopathological phenotypes (wang2021predicting; liu2025spatial).
@@
with reconstruction loss . Here, histology serves as the morphological manifestation of underlying molecular events, guiding omics-derived slots toward biologically meaningful alignment. Notably, because genomic profiling is substantially more costly than histopathology and often unavailable at test time (payne2018cost), this cross-modal reconstruction enhances robustness: the reconstruction head can impute genomic features directly from histology images, enabling inference even when omics data are missing.
Slots Interactions. To capture multimodal dependencies, we model slot interactions at both intra- and inter-modal levels—handled in parallel. Intra-modal interactions are captured using masked self-attention using only the top- most self-predictive slots within each modality. For inter-modal interactions, we use all slots without masking to fully capture inter-modal synergies, including those arising from slots that are originally less self-predictive within their own modality. Instead of relying on one-shot co-attention (chen2021multimodal; xu2023multimodal; jaume2024modeling), we design an iterative cross-attention scheme that follows the recurrent update in slot attention module. At each iteration, slots from one modality attend to those of the other: histology queries attend to omics keys/values, and vice versa (see Appendix B.4 for details). Finally, the output of three branches are concatenated and passed into a risk classification head for risk prediction. The computational complexity of slots interactions is . Overall, instead of the original , our model requires . Since in practice and the term is negligible, this expression can be upper-bounded by .

### v2

@@ Abstract
The integration of histology images and gene profiles has shown great promise for improving survival prediction in cancer. However, current approaches often struggle to model intra- and inter-modal interactions efficiently and effectively due to the high dimensionality and complexity of the inputs. A major challenge is capturing critical prognostic events that, though few, underlie the complexity of the observed inputs and largely determine patient outcomes. These events—manifested as high-level structural signals such as spatial histologic patterns or pathway co-activations—are typically sparse, patient-specific, and unannotated, making them inherently difficult to uncover. To address this, we propose SlotSPE, a slot-based framework for structural prognostic event modeling. Specifically, inspired by the principle of factorial coding, we compress each patient’s multimodal inputs into compact, modality-specific sets of mutually distinctive slots using slot attention. By leveraging these slot representations as encodings for prognostic events, our framework enables both efficient and effective modeling of complex intra- and inter-modal interactions, while also facilitating seamless incorporation of biological priors that enhance prognostic relevance. Extensive experiments on ten cancer benchmarks show that SlotSPE outperforms existing methods in 8 out of 10 cohorts, achieving an overall improvement of 2.9%. It remains robust under missing genomic data and delivers markedly improved interpretability through structured event decomposition. Code available at https://github.com/zylvemvet/SlotSPE .
@@ 1 Introduction
Cancer survival analysis aims to estimate patient mortality risk or survival time, serving as a key tool for personalized prognosis and precision oncology, as it helps quantify outcomes, identify prognostic factors, and guide individualized treatment decisions (Cox and Reid, 2004; Jenkins, 2005; Wiegrebe et al., 2024). Whole-slide images (WSIs) and genomics are two widely-used tools in cancer survival analysis, as they capture complementary prognostic signals: WSIs reveal how cancer manifests morphologically in tissue, while genomics uncovers its underlying molecular drivers. Their integration offers the opportunity for more accurate, robust, and biologically grounded survival prediction, which has in turn sparked growing interest in multimodal learning approaches for survival analysis (Chen et al., 2021b; Xu and Chen, 2023; Jaume et al., 2024; Zhou et al., 2025).
@@
Effective and efficient multimodal learning for survival analysis faces a fundamental challenge: although patient outcomes are often driven by a small number of critical events (Vogelstein et al., 2013; Hosseini et al., 2024), WSI and genomic signals are rather complex and high-dimensional. High-resolution WSIs can exceed pixels (Song et al., 2023) and genomics involve thousands of genes (Győrffy, 2021). These events manifest as meaningful structural associations within WSI and genomic signals—e.g., in histopathology, immune activity and tumor aggressiveness (Chan et al., 2024; Liang et al., 2025) are reflected in the density and proximity of the pixel organizations of tumor-infiltrating lymphocytes and cancer nests, while in genomics, the tumor-promoting effects emerge from the co-activation of PI3K/AKT and MAPK pathways (Glaviano et al., 2023). However, learning to uncover such event-related patterns structures from the input signals is difficult for machines: these high-level associations are typically sparse, patient-specific, and unannotated. Compounding the challenge, the enormous scale of input units—millions of WSI pixels and thousands of genomic features—creates massive self-interactions within each modality, further amplified by the need for cross-modal alignment, rendering direct modeling prohibitively expensive.
@@
Existing attempts, e.g., (Ilse et al., 2018; Shao et al., 2021; Chen et al., 2021b; Xu and Chen, 2023; Jaume et al., 2024), fall short of capturing high-level prognostic associations. The most relevant studies are prototype-based methods (Zhang et al., 2024; Song et al., 2024), which employ fixed prototype representations and therefore lack adaptability to individual patients. For example, PIBD (Zhang et al., 2024) approximates a prototypical distribution for each risk group via information bottleneck (Tishby et al., 2000; Alemi et al., 2016). Yet, these prototypes are static-once trained, the same set is applied to all patients. MMP (Song et al., 2024) takes a different route by adopting Gaussian mixture model (GMM) (Bishop and Nasrabadi, 2006) to define prototypes for each WSI. However, the image-wise prototypes are unlearnable (i.e., fixed) during training, making end-to-end optimization suboptimal.
@@
In this paper, we introduce SlotSPE, a slot-based framework for structural prognostic event modeling. Inspired by factorial coding (Schmidhuber, 1992; Higgins et al., 2017; Greff et al., 2020), our method compresses high-dimensional multimodal inputs into a compact set of semantic slots using a slot attention module (Locatello et al., 2020), where each slot corresponds to a latent prognostic event. Unlike fixed prototypes, these slots are dynamically instantiated for each patient, allowing the model to capture individualized prognostic patterns while preserving a concise representation. To reduce slot redundancy and encourage slots to capture distinct prognostic events, we introduce selective slot activation through a Mixture-of-Experts–style decoder (Hampshire and Waibel, 1989; Jacobs et al., 1991; Jordan and Jacobs, 1994; Shazeer et al., 2017), which activates only the most predictive slots for each patient.
Beyond compression, we further model cross-modal interactions by incorporating a biological prior: histopathological phenotypes often reflect underlying molecular events, revealing molecular-morphology mappings (Wang et al., 2021; Liu et al., 2025b). Guided by this insight, we design a cross-modal reconstruction task where omics-derived slots are encouraged to predict gene expression from histology image. This alignment not only enforces biologically meaningful interactions across modalities but also improves robustness in missing genomic data.
@@ 2 Related Works
Single Modality. The advent of deep learning marks a paradigm shift in cancer prognosis, where artificial neural networks (NN) can uncover prognostic patterns from radiological (Aerts et al., 2014; Lou et al., 2019; Xu et al., 2019; Yao et al., 2021), pathological (Ciresan et al., 2012; Cireşan et al., 2013), and molecular data (Chaudhary et al., 2018; Ching et al., 2018) that were previously inaccessible to conventional methods. In histopathology, whole-slide images (WSIs) under the multiple instance learning (MIL) (Dietterich et al., 1997; Maron and Lozano-Pérez, 1997) paradigm—where a slide is represented as a bag of patches with only slide- or patient-level survival labels—have become standard (Hou et al., 2016; Ilse et al., 2018; Shao et al., 2021; Lu et al., 2021). Recent MIL approaches address the challenge of abundant patches by using clustering (e.g. k-means) (Yao et al., 2020; Vu et al., 2023; Yang et al., 2024; Zhou et al., 2024), graph models (Li et al., 2018; Chen et al., 2021a), and hierarchical WSI representations (Shao et al., 2023). Meanwhile, genomic methods moved from simple MLP or shallow nets on high-dimensional gene profiles (Haykin, 1998; Klambauer et al., 2017) to pathway-based representations that group genes into biologically meaningful sets for better interpretability (Zhao et al., 2021; Hou et al., 2023; Wang et al., 2024). Building on these advances, our method preserves the MIL standard for histopathology, adopts pathway-level genomic features.
@@
Multiple Modality. Multimodal approaches, especially the integration of histopathology and genomics, have attracted increasing attention because combining morphological and molecular information often yields better prognostic performance than either modality alone (Chen et al., 2020; 2021b; Yuan et al., 2025). Typical strategies include late fusion (e.g., concatenation or Kronecker product) (Chen et al., 2020; Ding et al., 2023; Volinsky-Fremond et al., 2024; Zhao et al., 2025; Luo et al., 2025), and early fusion via co-attention mechanisms operating at the patch-gene level (Chen et al., 2021b; Xu and Chen, 2023; Zhou and Chen, 2023; Jaume et al., 2024; Yuan et al., 2025; Wang et al., 2025; Zhou et al., 2025). Since inputs are complex and high-dimensional, recent work has adopted adaptive token selection (Zhang et al., 2025) or compact latent representations by compressing WSI and pathway data via prototypes (Liu et al., 2025a; Song et al., 2024; Zhang et al., 2024). The former, however, lacks modeling of interpretable structural associations; the latter fails to address sparsity and patient-specific event variation, relying instead on fixed prototypes.
@@
Structural Event Modeling. A key challenge in multimodal survival analysis lies in handling the high-dimensional histology and omics inputs, which not only incurs substantial computational costs but also complicates modeling of intricate intra-/inter-modal interactions. An important insight is that patient outcomes are often driven by a small number of critical events (Vogelstein et al., 2013). This motivates approaches that aim to uncover the underlying event-related structures within complex inputs. For examples, PIBD (Zhang et al., 2024) and MMP (Song et al., 2024) attempt to derive structured representations from complex multimodal data. However, relying on fixed prototypes, these methods cannot dynamically capture the sparse and patient-specific structural events that often determine outcomes. From the perspective of machine learning, this challenge can be framed as a problem of factorial coding, a well-established paradigm concerned with decomposing and compressing high-dimensional data into disentangled latent factors (Schmidhuber, 1992; Tipping and Bishop, 1999; Higgins et al., 2017; Kim and Mnih, 2018). Building on this connection, our framework advances beyond static prototypes by introducing learnable slots through slot attention (Locatello et al., 2020), modeling prognostic events that are selectively activated per patient and aligned via biologically guided cross-modal reconstruction. This design enables a structured decomposition of survival-related events, binding multimodal prognostic signals to interpretable representations.
@@ 3 Method
Survival prediction is the task of estimating a patient risk of an event (e.g., death) given the pathological observations (e.g. WSIs and genomic) of the patient. In this paper, we tackle the problem of learning such hazard predictor from the multi-modal data. Formally, let the hazard prediction function be denoted as , representing the probability of death occurring at time t (i.e., survival up to time t). The model is trained using the negative log-likelihood (NLL) loss (Zadeh and Schmid, 2020), denoted as (see Appendix B.1 for details). Each patient sample is represented as , where indexes the patient, denotes the pathology data, denotes the genomic profile, , and is the censoring indicator ( for an observed event, for right-censoring). Operationally, following standard practices in computational pathology (Chen et al., 2021b; Jaume et al., 2024), WSI and genomic inputs are pre-processed by dedicated image and gene encoders and represented as bags of instances: and , where and denote the numbers of WSI patches and biological pathways, respectively, and is the feature dimension. When modeled with state-of-the-art transformers like prior works (Xu and Chen, 2023; Zhou and Chen, 2023; Jaume et al., 2024), the resulting intra- and inter-modal interactions scale quadratically as , creating a major challenge for both efficacy and efficiency.
@@ 3.1 Slot-based Information compression
Inspired by factorial coding (Schmidhuber, 1992; Higgins et al., 2017), which aims to disentangle data into independent latent factors, we perform structural decomposition for two bags of histology and genomic instances, i.e., and , using a slot attention module (Locatello et al., 2020). This allows us to compress the large input sets into much smaller slot-based representations, significantly reducing computational overhead. These slots are then decoded using a MoE-style decoder that selectively activates patient-specific slots, promoting both sparsity and competition among them.
Slot Attention. Slot attention (Locatello et al., 2020) learns to encode distinct structural patterns present in the input into a finite set of latent vectors, i.e., the “slots”. Since WSI and genomic data inherently contain massive redundancy (Zhang et al., 2024; Song et al., 2024), it is desirable to aggregate them into a smaller set of compact representations while preserving their semantic content. This makes slot attention particularly well-suited for our task, where the resulting compact representations can be interpreted as latent prognostic events modeled by slots.
We re-organize the input features for both modalities as matrices, . With a set of slots, organized as , initialized from some learnable distributions, the slot attention compresses into slots through iterations of cross-attention-based routing. At each iteration , the slots attend to the input to produce updated representations, which are then refined using a multilayer perceptron (MLP). Slot updates across iterations are applied recurrently using a recurrent neural network (RNNs) (Hochreiter and Schmidhuber, 1997; Cho et al., 2014) to preserve temporal consistency and iterative refinement. Mathematically, each slot is obtained by:
@@
Then the top- slots are selected using the Gumbel-Top- trick (Gumbel, 1954; Maddison et al., 2014; Kool et al., 2019) with Straight-Through Estimation (ST) (Jang et al., 2016), producing a differentiable -hot mask (see Appendix B.2 for details). The selected slots are reweighted by masking and renormalizing the softmax weights:
@@ 3.2 Multimodal Slots Interaction with Bilogical Guidance
Cross-modal Reconstruction using Biological Prior. To ensure effective cross-modal interaction, we introduce cross-modal reconstruction (Figure 2(b)) to uncover molecular-morphology mappings reflecting the fact that molecular events give rise to histopathological phenotypes (Wang et al., 2021; Liu et al., 2025b).
@@
with reconstruction loss . Here, histology serves as the morphological manifestation of underlying molecular events, guiding omics-derived slots toward biologically meaningful alignment. Notably, because genomic profiling is substantially more costly than histopathology and often unavailable at test time (Payne et al., 2018), this cross-modal reconstruction enhances robustness: the reconstruction head can impute genomic features directly from histology images, enabling inference even when omics data are missing.
Slots Interactions. To capture multimodal dependencies, we model slot interactions at both intra- and inter-modal levels—handled in parallel. Intra-modal interactions are captured using masked self-attention using only the top- most self-predictive slots within each modality. For inter-modal interactions, we use all slots without masking to fully capture inter-modal synergies, including those arising from slots that are originally less self-predictive within their own modality. Instead of relying on one-shot co-attention (Chen et al., 2021b; Xu and Chen, 2023; Jaume et al., 2024), we design an iterative cross-attention scheme that follows the recurrent update in slot attention module. At each iteration, slots from one modality attend to those of the other: histology queries attend to omics keys/values, and vice versa (see Appendix B.4 for details). Finally, the output of three branches are concatenated and passed into a risk classification head for risk prediction. The computational complexity of slots interactions is . Overall, instead of the original , our model requires . Since in practice and the term is negligible, this expression can be upper-bounded by .